How do you solve 2 (4/6)=5e
Answer by stanbon(75887)

How do you solve 2 (4/6)=5e
-------
2(2/3) = 5e
4/3 = 5e
---
e = (1/5)(4/3)
---
e = 4/15
------------------------
Cheers,
Stan H.
===================